### v1

# thanks: These authors contributed equallythanks: These authors contributed equally

Abstract
Non-Hermitian skin effects (NHSEs) have recently been investigated extensively at the single-particle level. When many-body interactions become dominant, novel non-Hermitian physical phenomena can emerge. In this work, we theoretically study NHSEs controlled by dissipation and interaction. We consider a 1D zigzag Bose-Hubbard lattice, subject to magnetic flux, staggered onsite single-particle loss, and uniform onsite two-particle loss. When the two-particle loss is small, two-body bound eigenstates (i.e., doublons) are all localized at the same boundary due to the interplay of the magnetic flux and staggered single-particle loss. While, for strong two-particle loss, the localization direction of doublons is unexpectedly reversed. This is attributed to the effective strong nonreciprocal hopping of doublons induced by the virtual second-order and third-order hopping processes of particle pairs with the combination of the magnetic flux, staggered single-particle loss, and strong two-particle loss. Moreover, a two-particle gain can induce the same skin-localization and reversal features of doublons, which can be utilized to dynamically observe the NHSE and its reversal of doublons controlled by interactions. Our results open up a new avenue for exploring novel non-Hermitian phenomena in many-body systems.
Introduction.—In the framework of non-Hermitian quantum mechanics Ashida et al. (2020), open quantum systems described by effective non-Hermitian Hamiltonians exhibit many striking physical phenomenon without Hermitian counterparts Peng et al. (2014a, b); Gao et al. (2015); Leykam et al. (2017); Xu et al. (2017); Gong et al. (2018); El-Ganainy et al. (2018); Yao and Wang (2018); Zhang et al. (2020); Yokomizo and Murakami (2019); Yao et al. (2018); Kunst et al. (2018); Liu et al. (2019); Song et al. (2019); Lee et al. (2019); Kawabata et al. (2019a); Lee et al. (2019); Zhang et al. (2018); Bliokh et al. (2019); Ge et al. (2019); Minganti et al. (2019); Zhao et al. (2019); Bliokh et al. (2019); Kawabata et al. (2019b); Borgnia et al. (2020); Minganti et al. (2020); Arkhipov et al. (2020); Okuma et al. (2020); Liu et al. (2021); Bergholtz et al. (2021); Li et al. (2022); Zhang et al. (2022a); Ren et al. (2022); Kawabata et al. (2023); Zhang et al. (2023); Li et al. (2023a); Leefmans et al. (2022); Okuma and Sato (2023); Li et al. (2024); Liu et al. (2023); Kawabata et al. (2024); Cai et al. (2024); Parto et al. (2023); Nie et al. (2023); Hu et al. (2024); Leefmans et al. (2024). One of the fascinating properties of non-Hermitian systems is the localization of bulk modes at the boundaries Yao and Wang (2018); Zhang et al. (2020); Yokomizo and Murakami (2019); Yao et al. (2018); Kunst et al. (2018); Liu et al. (2019); Song et al. (2019); Lee et al. (2019); Kawabata et al. (2019a); Lee et al. (2019), dubbed the non-Hermitian skin effect (NHSE). The NHSE originates from the intrinsic point-gap topology of complex eigenenergies Okuma et al. (2020). Recently, many efforts have been devoted to exploring unique consequences of NHSE without many-body interactions. For instance, it leads to the breakdown of conventional Bloch band theory Yao and Wang (2018); Yokomizo and Murakami (2019); Zhang et al. (2020), nonunitary scaling of non-Hermitian localization Kawabata and Ryu (2021), and entanglement phase transitions Kawabata et al. (2023).
When many-body interactions are introduced into non-Hermitian systems, unusual physical properties can emerge Ashida et al. (2017); Nakagawa et al. (2018); Yamamoto et al. (2019); Nakagawa et al. (2020); Yoshida and Hatsugai (2021); Zhang et al. (2022b); Dóra and Moca (2022); Kawabata et al. (2022); Liu et al. (2020); Yoshida and Hatsugai (2022); Shen and Lee (2022); Faugno and Ozawa (2022); Li et al. (2023b); Longhi (2023); Qin and Li (2024); Yoshida et al. (2023); Qin et al. (2024); Brighi and Nunnenkamp (2024); Kim et al. (2024). Especially, for interacting systems in the presence of nonreciprocal hopping, many-body interactions have the potential to induce NHSE Faugno and Ozawa (2022), occupation-dependent NHSE Qin and Li (2024), non-Hermitian Mott skin effect Yoshida et al. (2023), dynamical suppression of NHSE Qin et al. (2024), and so on. However, it is challenging requirement of nonreciprocal hopping in achieving the NHSE for many experimental platforms. A more operational approach is to utilize onsite loss or gain Li et al. (2020); Yi and Yang (2020); Li et al. (2022); Liang et al. (2022); Cai et al. (2024). Therefore, a natural open question arises: can the onsite loss or gain be utilized to induce and control NHSEs in the interacting system?
In this work, we study the skin-mode localization in dissipative and interacting systems without nonreciprocal hopping. Specifically, we consider a 1D zigzag lattice in the presence of a magnetic flux piercing through each triangular plaquette and a Bose-Hubbard interaction. For strong interaction, bound pairs, so-called doublon, are formed, where two bosons occupy the same site. By introducing the imbalanced onsite single-particle loss at the odd and even sites and uniform onsite two-particle loss, two branches of doublon bands form. For weak two-particle loss, the states from two doublon bands are all localized at the same boundary due to the interplay of the magnetic-flux-induced chiral transport and staggered single-particle loss. Most remarkably, the states from one branch of the doublon bands are unexpectedly reversely localized at the other boundary for strong two-particle loss. The underlying hidden physical mechanism is revealed by utilizing third-order quasidegenerate perturbation theory, where an effectively strong nonreciprocal hopping contributes from the virtual second-order and third-order hopping processes of particle pairs with the combination of the magnetic flux, staggered single-particle loss, and strong two-particle loss. We finally discuss the same reversal of the NHSE induced by adding two-particle gain. Moreover, with two-particle gain, the initial two-particle excitation can remain strongly bound during the dynamical evolution, which can be utilized to observe dynamical skin-mode localization and its reversal of doublons controlled by the interactions.
Model.—We consider a 1D zigzag bosonic chain in the presence of onsite many-body interactions, single-particle and two-particle dissipation, as shown in Fig. 1. In addition, a synthetic magnetic flux $\phi$ piercing through each triangular plaquette is introduced. The Bose-Hubard Hamiltonian is written as
[…]
$-\sum_{j=1}^{L}t\left(\hat{a}^{\dagger}_{j+1}\hat{a}_{j}+\text{H.c.}\right)-U\sum_{j=1}^{L}\hat{a}^{\dagger}_{j}\hat{a}^{\dagger}_{j}\hat{a}_{j}\hat{a}_{j},$ (1)
where $\hat{a}_{j}$ annihilates a boson at site $j$, $t$ and $t^{\prime}$ represent the nearest-neighbor and next-nearest-neighbor hopping strengths, $U$ denotes the many-body interacting strength, and $L$ is the number of lattice sites.
Figure 1: Tight-binding representation of a two-leg zigzag chain in the presence of onsite interaction $U$, imbalanced onsite single-particle loss (denoted by purple wavy lines) with decay rate $\gamma\pm\delta$ on two legs, and uniform onsite two-particle loss (denoted by yellow wavy lines) with decay rate $\kappa$. A synthetic magnetic flux $\phi$ piercing through each triangular plaquette is introduced. The black solid and red dashed lines represent the nearest-neighbor and next-nearest-neighbor hopping with strengths $t$ and $t^{\prime}$.
We consider this open quantum system, where the zigzag lattice is subjected to the imbalanced onsite single-particle loss between two legs, and uniform onsite two-particle inelastic collision (see Fig. 1). Under the Markovian approximation, the dissipative dynamics of the system is governed by the Lindblad master equation Scully and Zubairy (1997); Breuer and Petruccione (2007); Lidar (2020)
$\frac{d\hat{\rho}}{dt}=$ $-i\left[\hat{\mathcal{H}},~{}\hat{\rho}\right]+\sum_{j}\left[\gamma-(-1)^{j}\delta\right]\mathcal{D}[\hat{a}_{j}]\hat{\rho}$
$+\sum_{j}\kappa\mathcal{D}[\hat{a}_{j}\hat{a}_{j}]\hat{\rho},$ (2)
where $\hat{\rho}$ is the system density matrix, and the Lindblad superoperator $\mathcal{D}[\hat{\mathcal{L}}]\hat{\rho}=\hat{\mathcal{L}}\hat{\rho}\hat{\mathcal{L}}^{\dagger}-\{\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}},~{}\hat{\rho}\}/2$ represents bosonic dissipation. The imbalanced single-particle losses along the two legs are denoted by $\gamma\pm\delta$, and $\kappa$ is the two-particle decay rate.
Quantum-trajectory theory states that the dynamics in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) can be decomposed into nonunitary evolution and stochastic quantum jumps Meystre and Sargent (2007). By projecting out the quantum jumps by continuously monitoring the particle number Ashida et al. (2016); Nakagawa et al. (2020), the system is governed by the effective non-Hermitian Hamiltonian
^$\hat{\mathcal{H}}_{\text{nH}}=\hat{\mathcal{H}}-i\frac{\kappa}{2}\sum_{j=1}^{L}\hat{a}^{\dagger}_{j}\hat{a}^{\dagger}_{j}\hat{a}_{j}\hat{a}_{j}-\frac{i}{2}\sum_{j=1}^{L}\left[\gamma-(-1)^{j}\delta\right]\hat{n}_{j},$ (3)
where $\hat{n}_{j}=\hat{a}_{j}^{\dagger}\hat{a}_{j}$ is the density operator. The Hamiltonian ^$\hat{\mathcal{H}}_{\text{nH}}$ conserves the $U(1)$ symmetry, i.e., ^$\hat{\mathcal{H}}_{\text{nH}}$ commutes with the total particle number operator [^$[\hat{\mathcal{H}}_{\text{nH}},~{}\sum_{j}\hat{n}_{j}]=0$. Thus, we can study the arbitrary subspace of the particle number in the Fock basis. In this work, we consider the double-excitation subspace and assume $\phi=\pi/2$.
Figure 2: (a) Complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) in the presence of single-particle dissipation for $\kappa=0$. The discrete doublon bands on the far left are well separated from the continuum of scattering states. The zoom‐in view of the two doublon bands in (a) is shown in (b), where the magenta star and diamond denote the two branches of the doublon bands. The corresponding particle densities $\langle\hat{n}_{j}\rangle$ are shown in (c) and (d) under OBCs, where the red (blue) curves indicate the state distributions at the odd (even) sites. Other parameters are: $U=6$, $t=2$, $t^{\prime}=1$, $\gamma=4$, $\delta=2$, and $L=60$.
Single-Particle Dissipation Induced NHSE.—We first study the NHSE in the presence of only the single-particle dissipation with $\kappa=0$. We plot the complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under open boundary conditions (OBCs, blue dots) and periodic boundary conditions (PBCs, red dots) in Fig. 2(a). In the double-excitation subspace, the eigenspectrum is divided into continuous scattering states and discrete doublon bands. The scattering states consist of superpositions of two-particle states residing on different sites, while the doublon bands are composed of bound-particle pairs residing on the same site, with their eigenenergies well separated from the continuum of scattering states [see Fig. 2(a)]. In this work, we focus on two doublon bands, of which the zoom-in view on the far left of Fig. 2(a) is shown in Fig. 2(b). In the complex plane, each doublon band, indicated by the magenta star and diamond, forms a point gap (red dots) under PBCs. Each of these point gaps encircles the respective doublon band (blue dots) under OBCs, indicating usually the occurrence of NHSE. In order to demonstrate this indeed the case, we calculate the particle density as
[…]
where |$\ket{\psi_{m}}_{R}$ ($m=1,2,3,\cdots$) is the $m$th right eigenvector of the two-body non-Hermitian Hamiltonian with ^$\hat{\mathcal{H}}_{\text{nH}}\ket{\psi_{m}}_{R}=E_{m}\ket{\psi_{m}}_{R}$.
Figure 3: Complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) subjected to single- and two-particle dissipation for $\kappa=16$ and $\delta=2$. The zoom‐in view of the two doublon bands on the far left of (a) is shown in (b), where the magenta star and diamond denote two branches of doublon bands. The corresponding particle densities $\langle\hat{n}_{j}\rangle$ are shown in (c) and (d) under OBCs, where the red (blue) curves indicate the state distributions at the odd (even) sites. Other parameters are: $U=6$, $t=2$, $t^{\prime}=1$, $\gamma=4$, and $L=60$.
The site-resolved particle densities of two doublon bands are shown in Figs. 2(c,d), respectively, where red (blue) curves indicate state distributions at odd (even) sites. All the doublons are localized at the right boundary, indicating the NHSE. Moreover, the states from two branches of doublon bands occupy different legs of the zigzag lattice, where the states from the doublon band with larger (smaller) loss reside mainly the odd (even) sites [see Fig. 2(c,d)]. In order to reveal the hidden reasons for the appearance of the doublon skin-mode localization caused by the imbalanced onsite single-particle loss, we derive its effective Hamiltonian (see details in Supplementary Materials (SM) in Ref. Non ). The NHSE of doublon bands can be attributed to the interplay of the effective magnetic flux of doublons, threading through each triangular plaquette, and the renormalized imbalanced loss between its two legs. The non-zero flux leads to opposite particle transports along the odd and even sites Li et al. (2022); Cai et al. (2019). Due to the larger loss at the odd sites, the particle transport towards the right boundary are favored, and the backflow is suppressed. This leads to the skin-mode localization at the right boundary. Note that two isolated states of doublons, lying outside the point gaps [Fig. 2(b)], are localized at either the left or right boundary. There are two-particle Tamm-Shockley edge states Tamm (1932); Shockley (1939), resulting from the different renormalization of the system parameters at the edges and bulk (see details in Ref. Non ), which have no topological origin.
The NHSE originates from intrinsic non-Hermitian topology. In order to characterize the point-gap topology, we introduce a twist angle $\varphi$ ($\varphi\in[0,~{}2\pi]$) to PBC, and define the many-body winding number Zhang et al. (2022b); Kawabata et al. (2022); Yoshida and Hatsugai (2022)
$\mathcal{W}=\oint_{0}^{2\pi}\frac{d\varphi}{2\pi i}\frac{\partial}{\partial\varphi}\log\left[\det\left(\hat{\mathcal{H}}_{\text{nH}}-E_{\textrm{ref}}\right)\right],$ (5)
where $E_{\textrm{ref}}$ is the complex reference energy inside the point gap. The nonzero winding number with $\mathcal{W}=2$ for each doublon band indicates the intrinsic topological origin of doublons’ NHSE.
Two-Particle Dissipation Controlled NHSE.—We include the effects of two-particle loss, and assume the two-particle loss much larger than the single-particle one with $\kappa=16$ and $\delta=2$, and plot the complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) in Fig. 3(a). The zoom-in view of doublon bands is shown in Fig. 3(b). The eigenenergies of OBCs are enclosed by the point-gap loop, indicating the appearance of NHSE. In contrast to the case of single-particle dissipation, the two-particle loss reverses the localization direction of doublons, as shown in Fig. 3(c,d). The states from the doublon band with the larger loss occupy the odd sites and are still localized at the right boundary [see Fig. 3(c)]. However, the states from the other doublon band with the smaller loss are unexpectedly localized at the left boundary [see Fig. 3(d)], where they occupy the even sites. In addition, the winding number of the doublon band with the smaller loss becomes $\mathcal{W}=-2$, indicating also the reversal of NHSE. Furthermore, the localization direction of the doublon skin modes can be tuned by the interaction, single-particle loss or two-particle dissipation. For instance, by fixing the two-particle loss $\kappa$, but increasing the single-particle dissipation $\delta$, we observe once more the reversal of the localization direction of the doublon band Non .
In order to understand the underlying hidden mechanism of the unexpected reversal of the localization direction of the doublon skin modes, we derive the effective Hamiltonian ^$\hat{\mathcal{H}}_{\text{eff}}$ describing the doublon bands in the strong-interaction limit with $U\gg t,t^{\prime}$. In this case, the doublons are tightly bound, where their two-particle components reside on the same site. Moreover, the doublon states are energetically well-separated from two-particle scattering states. Therefore, we can split the Hamiltonian ^$\hat{\mathcal{H}}_{\text{nH}}$ into its dominant term, with ^$\hat{\mathcal{H}}_{0}=-(U+i\kappa/2)\sum_{j}\hat{n}_{j}(\hat{n}_{j}-1)-\frac{i}{2}\sum_{j}\left[\gamma-(-1)^{j}\delta\right]\hat{n}_{j}$, and treat the remaining terms as the perturbation. Based on third-order quasidegenerate perturbation theory Non , we obtain the effective Hamiltonian ^$\hat{\mathcal{H}}_{\text{eff}}$ of doublons as
^eff=subscript^$\hat{\mathcal{H}}_{\text{eff}}=$ $-t_{0}\sum_{j}\left(\hat{d}^{\dagger}_{j+1}\hat{d}_{j}+\textrm{H.c.}\right)+\sum_{j}\left(U_{1}\hat{n}_{2j-1}+U_{2}\hat{n}_{2j}\right)$
$-\sum_{j}\left(J_{\text{1,R}}\hat{d}^{\dagger}_{2j+1}\hat{d}_{2j-1}+J_{\text{1,L}}\hat{d}^{\dagger}_{2j-1}\hat{d}_{2j+1}\right)$
$-\sum_{j}\left(J_{\text{2,R}}\hat{d}^{\dagger}_{2j+2}\hat{d}_{2j}+J_{\text{2,L}}\hat{d}^{\dagger}_{2j}\hat{d}_{2j+2}\right),$ (6)
where $J_{\text{1,R}}=J(\delta,\phi)$, $J_{\text{1,L}}=J(\delta,-\phi)$, $J_{\text{2,R}}=J(-\delta,-\phi)$, $J_{\text{2,L}}=J(-\delta,\phi)$, $U_{1}=U(\delta)$, $U_{2}=U(-\delta)$, and
$t_{0}=$ $\left[\frac{2U}{4U^{2}+(\kappa+\delta)^{2}}+\frac{2U}{4U^{2}+(\kappa-\delta)^{2}}\right]t^{2}$
$-i\left(\frac{\kappa+\delta}{4U^{2}+(\kappa+\delta)^{2}}+\frac{\kappa-\delta}{4U^{2}+(\kappa-\delta)^{2}}\right)t^{2},$ (7)
$J(\delta,\phi)=$ $\frac{4t^{2}t^{\prime}}{(2U+i\kappa)(2U+i\kappa+i\delta)}e^{-i\phi}$
$\frac{2t^{2}t^{\prime}}{(2U+i\kappa+i\delta)^{2}}e^{-i\phi}+\frac{2t^{\prime 2}}{2U+i\kappa}e^{-2i\phi},$ (8)
$U(\delta)=$ $-\left[2U+\frac{8Ut^{2}}{4U^{2}+(\kappa+\delta)^{2}}+\frac{8Ut^{\prime 2}}{4U^{2}+\kappa^{2}}\right]$
$+i\left[\frac{4(\kappa+\delta)t^{2}}{4U^{2}+(\kappa+\delta)^{2}}+\frac{4\kappa t^{\prime 2}}{4U^{2}+\kappa^{2}}-\kappa-\gamma-\delta\right].$ (9)
According to the above doublon’s effective Hamiltonian in the 1D zigzag lattice, the magnetic flux piercing through each triangular plaquette [see Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)] and the imbalanced complex onsite potential $U(\pm\delta)$ [see Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)]) at odd and even sites are renormalized. Most remarkably, along each leg of the zigzag lattice, the virtual second-order hopping process of particle pairs contributes to an effective phase $2\phi$, and third-order process contributes to $\phi$ in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) (see details in Ref. Non ). Therefore, the effective next-nearest neighbor hopping of doublons becomes nonreciprocal, i.e., |$\absolutevalue{J_{i,\text{R}}}\neq\absolutevalue{J_{i,\text{L}}}$ | ($i=1,2$) in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects), due to the interplay of two-particle dissipation and magnetic flux. Moreover, the nonreciprocal-hopping direction can be different along two legs, determined by the interaction strength $U$, two-particle loss $\kappa$, and imbalanced single-particle loss $\delta$.
According to the above discussion, in the presence of two-particle dissipation, the skin-mode localization direction of doublons is determined by the competition between the effective nonreciprocal hopping, induced by the two-particle dissipation and magnetic flux, and the directional hopping due to the interplay of magnetic-flux induced chiral transport and imbalanced loss along two legs. This contrasts with the single-particle case, where the NHSE results from the interplay of the magnetic-flux-induced chiral transport and the imbalanced onsite loss along two legs of the zigzag lattice structure in Fig. 1. By fixing $U=6$ and $\delta=2$, we have |$\absolutevalue{J_{1,\text{R}}}>\absolutevalue{J_{1,\text{L}}}$ | and |$\absolutevalue{J_{2,\text{R}}}<\absolutevalue{J_{2,\text{L}}}$ | for $\kappa>\delta$. If the two-particle dissipation $\kappa$ is comparable with the interaction strength $U$, all the states from two doublon bands are localized at the right boundary (see discussions with $\kappa=7$ in Ref. Non ). This is mainly attributed to the directional hopping due to the interplay of the magnetic-flux-induced chiral transport and imbalanced loss along two legs of doublons [see Eqs. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) and (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)] in spite of the slightly larger hopping strength towards the left boundary along the even sites. However, as the two-particle dissipation strength $\kappa$ is further increased, the asymmetrical hopping towards the left boundary becomes much stronger than hopping towards the right one along the even sites. This finally reverses the localization direction of the states from the doublon band with the smaller loss for $\kappa\gg U$, where they are localized at the left boundary along the even sites [see Fig. 3(d)]. Furthermore, if $\delta$ is increased and becomes comparable with $\kappa$, we have |$\absolutevalue{J_{1,\text{R}}}>\absolutevalue{J_{1,\text{L}}}$ | and |$\absolutevalue{J_{2,\text{R}}}>\absolutevalue{J_{2,\text{L}}}$ |. This causes once more the reversal of NHSE along the even sites [see Fig. 3(h)].
Figure 4: Complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nHG}}$ under OBCs (blue dots) and PBCs (red dots) subject to the two-particle gain (a) for $\kappa=-0.5$ and (e) for $\kappa=-2$. The magenta star and diamond denote the two doublon bands on the far left of both (a) and (e). The corresponding particle densities $\langle\hat{n}_{j}\rangle$ are shown in (b,c) and (f,g) under OBCs, where the red (blue) curves indicate state distributions at the odd (even) sites. Evolution dynamics of the doublon density $\langle\hat{n}_{j}^{d}\rangle$ with a double-occupied state initially prepared at the odd site (d) for $\kappa=-0.5$, and (h) for $\kappa=-2$. Other parameters are: $U=0.5$, $t=0.2$, $t^{\prime}=0.1$, $\gamma=1$, and $\delta=0.5$.
Two-Particle Pumping and Dynamics.—We now proceed to study the evolution dynamics of doublons. When the system is subjected to both single-particle and two-particle dissipation, the state decays during the dynamical evolution, and collapses into the system’s eigenstate with the largest imaginary part of its eigenenergy after a long-time evolution. For the double-excitation subspace, the eigenenergies of scattering states have the larger imaginary parts than those of doublons (see Fig. 3). Therefore, the system eventually evolves into the scattering state in spite of the initial doublon excitation, leading to the doublon dissociation Longhi (2023).
According to the above effective Hamiltonian of doublons in the strong-interaction limit, we infer that the two-particle gain with $\kappa<0$ plays the same role in determining the localization direction of doublons as the two-particle loss with $\kappa>0$. Therefore, in order to dynamically observe the NHSE reversal of doublons controlled by the interaction, we study the NHSE by introducing the two-particle gain along the odd sites via incoherent two-particle pumping. The effective non-Hermitian Hamiltonian becomes
^nHG=subscript^$\hat{\mathcal{H}}_{\text{nHG}}=~{}$$\hat{\mathcal{H}}-i\frac{\kappa}{2}\sum_{j=1}^{L/2}\hat{a}^{\dagger}_{2j-1}\hat{a}^{\dagger}_{2j-1}\hat{a}_{2j-1}\hat{a}_{2j-1}$
$-\frac{i}{2}\sum_{j=1}^{L}\left[\gamma-(-1)^{j}\delta\right]\hat{n}_{j},$ (10)
where $\kappa<0$. By setting $U=0.5$, $\gamma=1$ and $\delta=0.5$, all the doublons are localized at the right boundary for small two-particle gain with $\kappa=-0.5$ (see Fig. 4(a-c) and also Ref. Non ) due to the interplay of the magnetic-flux-induced chiral transport and the larger onsite loss along the odd sites. For the large two-particle gain with $\kappa=-2$, the effective loss (it is equivalent to add a large background loss) at odd sites becomes smaller than that at even sites. However, not all the doublon states are localized towards the left boundary, and the states from one doublon band are still unexpectedly localized at the right boundary along the even sites [see Fig. 4(e-g)].
The dynamical evolution of the system is calculated according to
[…]
where the initial state is chosen as the double-occupied state (i.e., a doublon) at the site $j$ with |$\ket{\psi_{0}}=1/\sqrt{2}\hat{a}_{j}^{\dagger}\hat{a}_{j}^{\dagger}\ket{0}$. Then, we calculate the time-resolved doublon density $\langle\hat{n}_{j}^{d}\rangle=\bra{\psi(t)}\hat{a}_{j}^{\dagger}\hat{a}_{j}^{\dagger}\hat{a}_{j}\hat{a}_{j}\ket{\psi(t)}/2$.
As shown in Fig. 4(d,h), in the presence the two-particle gain $\kappa$, the onsite two-particle excitation is strongly bound during the dynamical evolution for the double-occupied state initially prepared at the odd site. For small |$\absolutevalue{\kappa}$ | with $\kappa=-0.5$, as expected, the doublon transports towards the right boundary [see Fig. 4(d)]. While, for the large two-particle gain with $\kappa=-2$, the evolution direction is reversed [see Fig. 4(h)], indicating the reversal of NHSE controlled by the two-particle gain.
Conclusion.—We investigate the loss or gain controlled NHSEs, with the combination of the Bose-Hubbard interaction, in a 1D zigzag lattice subject to a magnetic flux threading through each triangular plaquette. By introducing imbalanced onsite single-particle loss on the odd and even sites, we realize all the states, from two doublon bands, localized at the same boundary due to the interplay of the magnetic-flux-induced chiral transport and staggered single-particle loss. Imposing the uniform onsite two-particle loss or gain produces an unexpected reversal of the localization direction of one doublon band due to the effectively strong nonreciprocal hopping contributing from the virtual second-order and third-order hopping processes of particle pairs. In experiments, our predictions could be tested experimentally using ultracold atoms Liang et al. (2022); Ren et al. (2022); Cai et al. (2019) and dissipative photonic resonator Leefmans et al. (2024), and the dynamical skin-mode localization and its reversal can be observed via introducing incoherent two-particle pumping. There can also be simulated in the electrical circuits by mapping the 1D two-particle excitation into a 2D lattice Olekhno et al. (2020). In the future, it is interesting to explore the NHSEs controlled by many-body interaction and loss beyond double-excitation subspace.
Acknowledgements.
T.L. acknowledges the support from National Natural Science Foundation of China (Grant No. 12274142), Introduced Innovative Team Project of Guangdong Pearl River Talents Program (Grant No. 2021ZT09Z109), the Fundamental Research Funds for the Central Universities (Grant No. 2023ZYGXZR020), and the Startup Grant of South China University of Technology (Grant No. 20210012). F.N. is supported in part by: Nippon Telegraph and Telephone Corporation (NTT) Research, the Japan Science and Technology Agency (JST)[via the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&\&&D Grant Number JPMJMS2061], the Asian Office of Aerospace Research and Development (AOARD) (via Grant No. FA2386-20-1-4069), and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074).